22cmcrawfkrd:

A regular heptagon has an apothem of approximately 7.3 yd and a perimeter of approximately 49.0 yds. Based on these measurements, what is the area of the heptagon?

AZ:

We have a heptagon (which means it's a polygon with 7 sides) The apothem is the distance from the center to the midpoint of a side. This is an example shown on a hexagon (six sided figure)

1 attachment
AZ:

To calculate the area of a regular heptagon, we can just add up the area of the 7 triangles Area of triangle = 1/2 * base * height We can calculate the base from the perimeter. The perimeter of the heptagon is the sum of all the sides. To find the length of each side, just divide the perimeter by 7 49 /7 = ?? That's the base of the triangle The apothem is the height of the triangle which is 7.3 yards

1 attachment
AZ:

So to find the area of the heptagon, find the area of one triangle area = 1/2 * base * height plug in the numbers you got from my previous reply the height is 7.3 the base is going to be 49/7 which is?? and so that's the area of one triangle The heptagon has 7 triangles so multiply that area by 7. That will be your final answer.